Quantity supplied has decreased Demand has decreased. O Demand has increased Supply has increased. Supply has decreased. QUESTION 6 Assume that the market for Good X is defined as follows: Q0 64-16P and Qs 16P- 8. What is the equilibrium price and output? s2.25, 28 $2.16 s2.75, 36 $4,48 $4,16 QUESTION 7 Assume that the market for Good X is defined as follows: QD-64-16P and QS-16P-8. What is consumer surplus in this market? O $40.50 $24.50 QUESTIONS Assume that the market for Good X is defined as follows: OD·64-16P and QS-16. 8. If the government restricts output in this market to 16 Both questions 6 and 7
Answer #1

a) At the equilibrium the price and quantity will be the same i.e. the demand and the supply are equal. So, we can write it as Qd=Qs.

64 - 16P = 16p - 8

= 64 + 8 = 16P + 16P

= 72 = 32P

P = 2.25. At the equilibrium the price will be 2.25. quantity will be Qd = 64 - 16(2.25) = 64 - 36 = 28.

Equilibrium price is 2.25 and quantity is 28. The answer is "A".

b) Price when the demand is zero = 64 -16P.

= 64 = -16P = P = 64 / 16 P = 4. when the price is 4 the demand is zero.

At the equilibrium the price is 2.25 and quantity is 28.

Consumer surplus = 1/2 ( 4 - 2.25 x 28 )

= 24.50. the answer is "C".